### v1

# TreeFive.bib

Abstract.
We prove that a tree with a vertex of degree at least five must be missing a connected partition of some type and therefore its chromatic symmetric function cannot be $e$-positive. We prove that this also holds for a tree with a vertex of degree four as long as it is not adjacent to any leaf. This brings us very close to the conjecture by Dahlberg, She, and van Willigenburg of non-$e$-positivity for all trees with a vertex of degree at least four. We also prove that spiders with four legs cannot have an $e$-positive chromatic symmetric function.
Key words and phrases:
chromatic symmetric function, connected partition, cut vertex, e-positive, proper colouring, spider graph, Stanley–Stembridge conjecture, tree isomorphism conjecture
2020 Mathematics Subject Classification:
Primary 05E05; Secondary 05C05, 05C15
Department of Mathematics, MIT
1. Introduction
When Stanley [chromsym] defined the chromatic symmetric function $X_{G}$ of a graph $G$, he posed two monumental problems that have garnered tremendous interest. The first is the Stanley–Stembridge conjecture [stanstem], which asserts that $X_{G}$ is $e$-positive whenever $G$ is the incomparability graph of a $(\bm{3}+\bm{1})$-free poset. Many authors proved $e$-positivity for particular subclasses of such graphs [csfvertex, chrombounce3, chromstoe, chromqrook, trilad, chromnsym, qforesttriples] and studied connections to Hessenberg varieties [chromhesssplitting, dothessenberg, hessenberghopf, stanstemhess, chromquasihessenberg] and Hecke algebras [chromhecke, chromcharacters]. The second is the tree isomorphism conjecture, which asserts that trees $T_{1}$ and $T_{2}$ can have $X_{T_{1}}=X_{T_{2}}$ only if they are isomorphic. This is intriguing because all trees with $n$ vertices have the same chromatic polynomial. Many authors proved that the chromatic symmetric function determines certain invariants and therefore distinguishes particular subclasses of trees [propcatchrom, chromtreepolyinv, markedgraphs, propqcatchrom, algochromtrees, catchromiso, distinguishingtrees, chromequal].
Dahlberg, She, and van Willigenburg [chromspostrees] considered a hybrid of these two questions by asking which trees $T$ have that $X_{T}$ is $e$-positive. Many authors [eposclawcon, spider4m22m1, spiderbroom, spiderepos] proved $e$-positivity or non-$e$-positivity for particular subclasses of trees. Dahlberg, She, and van Willigenburg checked the following conjecture for all trees with at most $12$ vertices.
Conjecture 1.1.
[chromspostrees, Conjecture 6.1] If $T$ is a tree with a vertex of degree $d\geq 4$, then $X_{T}$ is not $e$-positive.
A powerful tool for studying $e$-positivity is the following necessary condition by Wolfgang.
Proposition 1.2.
[conparts, Proposition 1.3.3] A connected graph $G$ must have a connected partition of every type in order for $X_{G}$ to be $e$-positive.
Note that this condition is not sufficient because for example, for every $m\geq 1$ the spider graph $G=S(2m+2,1,1)$ has a connected partition of every type but $X_{G}$ is not $e$-positive. Dahlberg, She, and van Willigenburg proved the following result.
Theorem 1.3.
[chromspostrees, Theorem 4.1] If $G$ is an $n$-vertex connected graph with a cut vertex whose deletion produces a graph with $d\geq\log_{2}n+1$ connected components, then $G$ is missing a connected partition of some type and therefore $X_{G}$ is not $e$-positive.
In particular, if $T$ is an $n$-vertex tree with a vertex of degree $d\geq\log_{2}n+1$, then $X_{T}$ is not $e$-positive. Zheng strengthened this result considerably.
Theorem 1.4.
[spiderepos, Theorem 3.13] If $G$ is a connected graph with a cut vertex whose deletion produces a graph with at least $6$ connected components, then $G$ is missing a connected partition of some type.
Our main result lowers this bound further.
Theorem 1.5.
Let $G$ be a connected graph with a cut vertex whose deletion produces a graph with either at least $5$ connected components, or $4$ connected components that each have at least two vertices. Then $G$ is missing a connected partition of some type.
In particular, if $T$ is a tree with a vertex of degree at least $5$, or with a vertex of degree $4$ that is not adjacent to any leaf, then $T$ is missing a connected partition of some type and therefore $X_{T}$ is not $e$-positive.
Note that there exists a tree $T$ with a vertex of degree $4$ that is adjacent to a leaf, and which has a connected partition of every type. Therefore, in this sense, Theorem 1.5 cannot be improved. In fact, we will show in Proposition 3.19 that for every $m\geq 1$ the spider graph $S(6m,6m-2,1,1)$ has a connected partition of every type, so there are infinitely many such examples. This means that to prove non-$e$-positivity of trees with a vertex of degree $4$ in general, we need to know more about the tree to calculate particular coefficients. In the specific case of spiders, we will prove that Conjecture 1.1 holds.
Theorem 1.6.
If $S$ is a spider with four legs, then $X_{S}$ is not $e$-positive.
Note that there exist infinite families of spiders with three legs, such as $S(m+1,m,1)$ for $m\geq 1$, that have $e$-positive chromatic symmetric function [chromspostrees, Theorem 3.2].
In the next section, we define chromatic symmetric functions, connected partitions, and spider graphs. In Section 3, we prove Theorem 1.5 by analyzing partial sums of rearrangements of partitions.
2. Background
A composition is a finite sequence of positive integers $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$, which we call the parts of $\alpha$. We denote by
$\text{set}(\alpha)=\{\alpha_{1},\alpha_{1}+\alpha_{2},\ldots,\alpha_{1}+\cdots+\alpha_{\ell-2},\alpha_{1}+\cdots+\alpha_{\ell-1}\}$
the set of (proper) partial sums of $\alpha$. A partition is a composition $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$ that also satisfies $\lambda_{1}\geq\cdots\geq\lambda_{\ell}$. We write $\lambda\vdash n$ if $\lambda_{1}+\cdots+\lambda_{\ell}=n$. We say that $\alpha$ is a rearrangement of $\lambda$ if every part occurs the same number of times in $\alpha$ and in $\lambda$.
Example 2.1.
The composition $\alpha=(3,4,2,4)$ has $\text{set}(\alpha)=\{3,7,9\}$ and is a rearrangement of the partition $\lambda=(4,4,3,2)\vdash 13$.
Let $G=(V,E)$ be a graph with $n$ vertices. A proper colouring of $G$ is a function $\kappa:V\to\{1,2,3,\ldots\}$ such that whenever vertices $u,v\in V$ are adjacent, we must have $\kappa(u)\neq\kappa(v)$. The chromatic symmetric function of $G$, defined by Stanley [chromsym, Definition 2.1] is the formal power series in infinitely many commuting variables $(x_{1},x_{2},x_{3},\ldots)$ given by
$X_{G}=\sum_{\kappa\text{ proper colouring }}\prod_{v\in V}x_{\kappa(v)}.$
We can write $X_{G}$ uniquely as a linear combination of elementary symmetric functions $e_{\lambda}$, which are indexed by partitions $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})\vdash n$ and defined by
$e_{\lambda}=e_{\lambda_{1}}\cdots e_{\lambda_{\ell}},\text{ where }e_{k}=\sum_{i_{1}\leq\cdots\leq i_{k}}x_{i_{1}}\cdots x_{i_{k}}.$
We say that $X_{G}$ is $e$-positive if all of the coefficients $c_{\lambda}$ in the expansion $X_{G}=\sum_{\lambda\vdash n}c_{\lambda}e_{\lambda}$ are nonnegative. It is also common to say that $G$ itself is $e$-positive. A connected partition of $G$ of type $\lambda\vdash n$ is a collection $\mathcal{S}=\{S_{1},\ldots,S_{\ell}\}$ of pairwise disjoint subsets of $V$ such that
  • $\circ$
    $\bigcup_{i=1}^{\ell}S_{i}=V$,
  • $\circ$
    the induced subgraph $G[S_{i}]$ is connected for every $1\leq i\leq\ell$, and
  • $\circ$
    $|S_{i}|=\lambda_{i}$ for every $1\leq i\leq\ell$.
Proposition 2.2.
[conparts, Proposition 1.3.3] Let $G$ be a connected $e$-positive graph with $n$ vertices. Then for every partition $\lambda\vdash n$, there exists a connected partition of $G$ of type $\lambda$.
A tree is a connected acyclic graph. A leaf of a tree is a vertex of degree one. A cut vertex of a connected graph $G$ is a vertex whose deletion results in a disconnected graph. For a partition $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$, we denote by $S(\lambda)$ the graph consisting of paths of lengths $\lambda_{1},\ldots,\lambda_{\ell}$, called legs, joined at a common vertex. Graphs of the form $S(\lambda)$ are called spiders.
Example 2.3.
Figure 1 shows some spiders and their chromatic symmetric functions. The spider $S(1,1,1)$ is missing a connected partition of type $(2,2)$ and is indeed not $e$-positive. The spider $S(3,2,1)$ has a connected partition of type $\lambda$ for every $\lambda\vdash 7$ and is $e$-positive. The spider $S(4,1,1)$ has a connected partition of type $\lambda$ for every $\lambda\vdash 7$ but is not $e$-positive, showing that Wolfgang’s necessary condition for $e$-positivity is not sufficient.
Figure 1. Some spiders and their chromatic symmetric functions
$S(1,1,1)$         $S(3,2,1)$         $S(4,1,1)$
$X_{S(1,1,1)}$ $=e_{211}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-2e_{22}}+5e_{31}+4e_{4}$
$X_{S(3,2,1)}$ $=e_{2221}+2e_{3211}+5e_{322}+4e_{331}+12e_{421}+5e_{43}+4e_{511}+13e_{52}+11e_{61}+7e_{7}$
$X_{S(4,1,1)}$ $=e_{2221}+4e_{3211}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-3e_{322}}+10e_{331}+10e_{421}+17e_{43}+4e_{511}+3e_{52}+11e_{61}+7e_{7}$
3. Proof of main result
We first fix notation that we will use throughout this section. Let $G$ be an $n$-vertex connected graph with a cut vertex $v$ whose deletion produces a graph with at least $3$ connected components, and denote their sizes
$a\geq b\geq c_{1}\geq\cdots\geq c_{k}\geq 1,\text{ where }k\geq 1.$
See Figure 2. Let $c=c_{1}+\cdots+c_{k}$. We will always assume that $c\geq 2$. Note that
$n=a+b+c+1\geq 2b+c+1.$
We now state our main result using this notation.
Theorem 3.1.
Suppose that $c\geq 2$. If any of the following hold, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
  1. (1)
    We have $b\leq 2c-2$.
  2. (2)
    We have $c\geq c_{1}+1$ and $b=2c-1$.
  3. (3)
    We have $2c\leq b\leq\frac{c^{2}}{2}$.
  4. (4)
    We have $c\geq c_{1}+2$ and $b\geq\frac{c^{2}}{2}$.
In particular, if $c\geq c_{1}+2$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Remark 3.2.
The condition that $c\geq c_{1}+2$ is precisely that the deletion of $v$ produces either at least $5$ connected components, or $4$ connected components that each have at least two vertices, so Theorem 3.1 implies Theorem 1.5. We cannot remove this hypothesis because the spider graph $S(6,4,1,1)$ has a connected partition of every type $\lambda\vdash 13$.
Remark 3.3.
The condition that $c\geq c_{1}+1$ is precisely that the deletion of $v$ produces at least $4$ connected components. We cannot remove this hypothesis because the spider graph $S(5,3,2)$ has a connected partition of every type $\lambda\vdash 11$ and is in fact $e$-positive.
Remark 3.4.
We cannot remove the hypothesis that $c\geq 2$ because the family of spider graphs $S(m+1,m,1)$ all have a connected partition of every type $\lambda\vdash 2m+3$ and are in fact $e$-positive.
We begin by describing another way to think about connected partitions.
Lemma 3.5.
Let $\lambda\vdash n$ be a partition with all parts at least $(c_{1}+1)$. If every rearrangement $\alpha$ of $\lambda$ has a partial sum $s\in\text{set}(\alpha)$ that lies in the interval $I=\{b+1,b+2,\ldots,b+c-1,b+c\}$, then $G$ has no connected partition of type $\lambda$.
Proof.
Suppose that $G$ has a connected partition $\mathcal{S}$ of type $\lambda$. We will find a rearrangement $\alpha$ of $\lambda$ that has no partial sum in $I$. Let $A,B,C_{1},\ldots,C_{k}$ be the connected components of $G\setminus v$ whose sizes are respectively $a\geq b\geq c_{1}\geq\cdots\geq c_{k}$ and let $S\in\mathcal{S}$ be the subset with $v\in S$. If any $C_{i}\nsubseteq S$, then letting $u\in C_{i}\setminus S$, the subset $S^{\prime}\in\mathcal{S}$ with $u\in S^{\prime}$ would have $|S^{\prime}|\leq c_{i}\leq c_{1}$, which is impossible because every part of $\lambda$ is at least $(c_{1}+1)$. So every $C_{i}\subseteq S$. This means that $\mathcal{S}$ is of the form $\mathcal{S}=\{A_{1},\ldots,A_{i},S,B_{1},\ldots,B_{j}\}$ where $A_{i^{\prime}}\subseteq A$ for $1\leq i^{\prime}\leq i$ and $B_{j^{\prime}}\subseteq B$ for $1\leq j^{\prime}\leq j$. We now take $\alpha$ to be the composition
$\alpha=(|B_{1}|,\ldots,|B_{j}|,|S|,|A_{1}|,\ldots,|A_{i}|).$
By definition, $\alpha$ is a rearrangement of $\lambda$, and we have
$\alpha_{1}+\cdots+\alpha_{j}$ $=|B_{1}|+\cdots+|B_{j}|\leq|B|=b\text{ and }$
$\alpha_{1}+\cdots+\alpha_{j+1}$ $=|B_{1}|+\cdots+|B_{j}|+|S|\geq|B|+|C_{1}|+\cdots+|C_{k}|+1=b+c+1.$
Therefore, no partial sum of $\alpha$ lies in the interval $I=\{b+1,\ldots,b+c\}$. ∎
Figure 2. A graph $G$ with a cut vertex $v$ joining at least $3$ connected components
[…]
We can now prove Part 4 of Theorem 3.1, the case where $b\geq\frac{c^{2}}{2}$.
Proposition 3.6.
If $c\geq c_{1}+2$ and $b\geq\frac{c^{2}}{2}$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
The idea is to take a partition $\lambda\vdash n$ with all parts equal to $(c-1)$ or $c$. The well-known Frobenius coin problem tells us that every integer $n\geq(c-1)(c-2)$ can be written in the form $n=a_{1}c+a_{2}(c-1)$ for some integers $a_{1},a_{2}\geq 0$. Because $b\geq\frac{c^{2}}{2}$, we have
$n\geq 2b+c+1\geq c^{2}+c+1\geq(c-1)(c-2),$
so we can write $n=a_{1}c+a_{2}(c-1)$ and take $\lambda$ to be $a_{1}$ $c$’s followed by $a_{2}$ $(c-1)$’s. All parts of $\lambda$ are at least $c-1\geq c_{1}+1$. Let $\alpha$ be a rearrangement of $\lambda$. Because $c\geq c_{1}+2\geq 3$, we have $\alpha_{1}\leq c\leq\frac{c^{2}}{2}\leq b$, so there exists some maximal $i$ with $\alpha_{1}+\cdots+\alpha_{i}\leq b$, and then
$b+1\leq\alpha_{1}+\cdots+\alpha_{i}+\alpha_{i+1}\leq b+\alpha_{i+1}\leq b+c.$
Therefore the result follows from Lemma 3.5. ∎
We now describe another way to ensure that all rearrangements have a partial sum in $I$.
Lemma 3.7.
Let $q$ be a positive integer and let $J=\{x,x+1,\ldots,y-1,y\}$, where
(3.1) $x=\left\lceil\frac{b+1}{q}\right\rceil\text{ and }y=\left\lfloor\frac{b+c}{q}\right\rfloor.$
Suppose that $x\geq c_{1}+1$. If $\lambda\vdash n$ is a partition with all parts in the interval $J$, then $G$ has no connected partition of type $\lambda$.
Proof.
All parts of $\lambda$ are at least $x\geq c_{1}+1$ and every rearrangement $\alpha$ of $\lambda$ has
$b+1\leq qx\leq\alpha_{1}+\cdots+\alpha_{q}\leq qy\leq b+c,$
so the result follows from Lemma 3.5. ∎
We now generalize the result of the Frobenius coin problem by seeing which numbers can be written as a sum of numbers in an interval $J$.
Lemma 3.8.
Let $x$ and $y$ be positive integers with $x<y$. Then there exists a partition $\lambda\vdash n$ with all parts in the interval $J=\{x,\ldots,y\}$ as long as
(3.2) $n\geq\left\lceil\frac{x-1}{y-x}\right\rceil x.$
Proof.
For a positive integer $t$, let $tJ=\{tx,tx+1,\ldots,ty-1,ty\}$. First note that every integer $n\in tJ$ has a partition $\lambda\vdash n$ consisting of $t$ integers in $J$. Indeed, for $n=ty$ we can take $\lambda=(y,\ldots,y)$, and for $tx\leq n\leq ty-1$ we can write $n=tq+r$ for some integers $x\leq q\leq y-1$ and $10\leq r\leq t-$ 1, and take $\lambda$ to be $r$ $(q+1)$’s followed by $(t-r)$ $q$’s.
We now see that as $t$ increases, the intervals $tJ$ become closer to each other. Specifically, for $t\geq\left\lceil\frac{x-1}{y-x}\right\rceil\geq\frac{x-1}{y-x}$, we have $ty+1\geq(t+1)x$, which means that as soon as we exit the interval $tJ$, we are already in the interval $(t+1)J$. This means that every integer $n$ satisfying (3.2) is in some interval $tJ$ and has a partition $\lambda\vdash n$ with all parts in $J$. ∎
Example 3.9.
Let $x=7$ and $y=9$, so that $J=\{7,8,9\}$ and we can write the numbers
$\{\underbrace{7,8,9}_{J}\}\cup\{\underbrace{14,15,16,17,18}_{2J}\}\cup\{\underbrace{21,22,23,24,25,26,27}_{3J}\}\cup\{\underbrace{28,29,30,31,32,33,34,35,36}_{4J}\}\cup\cdots$
as a sum of numbers in $J$. We see that once $t\geq\left\lceil\frac{x-1}{y-x}\right\rceil=3$, we no longer have a gap between the intervals $tJ$ and $(t+1)J$, so we are able to express any integer $n\geq 3x=21$.
Proposition 3.10.
If $n\geq\left\lceil\frac{b}{c-1}\right\rceil(b+1)$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
By Lemma 3.8, there exists a partition $\lambda\vdash n$ with all parts in the interval $J=I=\{b+1,\ldots,b+c\}$. All parts of $\lambda$ are at least $b+1\geq c_{1}+1$ and trivially every rearrangement $\alpha$ of $\lambda$ has the partial sum $\alpha_{1}\in\{b+1,\ldots,b+c\}$, so the result follows from Lemma 3.5. ∎
We can now prove Parts 1 and 2 of Theorem 3.1.
Proposition 3.11.
If $c\geq 2$ and $b\leq 2c-2$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
If $b\leq 2c-2$, then we have
$\left\lceil\frac{b}{c-1}\right\rceil(b+1)\leq 2(b+1)\leq 2b+c+1\leq n,$
so the result follows from Proposition 3.10. ∎
Proposition 3.12.
If $c\geq c_{1}+1$ and $b=2c-1$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
First note that if $c=2$, then $b=3$, $I=\{4,5\}$, and taking $\lambda=(2,\ldots,2)$ if $n$ is even or $\lambda=(3,2,\ldots,2)$ if $n$ is odd, every rearrangement $\alpha$ has $\alpha_{1}+\alpha_{2}=4$ or $5$ and we are done by Lemma 3.5. Now suppose that $c\geq 3$. Let $x=\left\lceil\frac{b+1}{2}\right\rceil=c\geq c_{1}+1$ and $y=\left\lfloor\frac{b+c}{2}\right\rfloor=\left\lfloor\frac{3c-1}{2}\right\rfloor$. By Lemma 3.8, we can find a partition $\lambda\vdash n$ with all parts in $J=\{x,\ldots,y\}$ because
$\left\lceil\frac{x-1}{y-x}\right\rceil x\leq\left\lceil\frac{c-1}{\frac{c-2}{2}}\right\rceil c=\left\lceil 2+\frac{2}{c-2}\right\rceil c\leq 4c\leq 2b+c+1\leq n.$
By Lemma 3.7 with $q=2$, $G$ has no connected partition of type $\lambda$. ∎
Our last case to prove is Part 3, where $2c\leq b\leq\frac{c^{2}}{2}$. We first check several cases by computer.
Proposition 3.13.
If $2\leq c\leq 40$ and $2c\leq b\leq\frac{c^{2}}{2}$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
By Proposition 3.10, we may assume that $2b+c+1\leq n\leq\left\lceil\frac{b}{c-1}\right\rceil(b+1)$, so there are finitely many triples $(b,c,n)$ to consider. For each one, we checked by computer that there exists a positive integer $q$ and a partition $\lambda\vdash n$ with all parts in the corresponding interval $J=\{x,\ldots,y\}$, where $x\geq c+1\geq c_{1}+1$, so the result follows from Lemma 3.7. ∎
Lemma 3.14.
Suppose that we can find a positive integer $q$ such that the integers $x$ and $y$ as in (3.1) satisfy $x\geq c+1$, $x<y$, and
(3.3) $\left\lceil\frac{x-1}{y-x}\right\rceil x\leq 2b+c+1.$
Then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
This follows directly from Lemma 3.7 and Lemma 3.8. ∎
Proposition 3.15.
If $41\leq c\leq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
There are finitely many pairs $(b,c)$ to consider. For each one, we checked by computer that there exists a positive integer $q$ such that $x\geq c+1$, $x<y$, and (3.3) holds. ∎
We now come to the most technical part of the proof, where given any $c\geq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$, we will find a suitable integer $q$. We first make a remark about this inequality.
Remark 3.16.
If we ignore the floors, ceilings, and ones, we have a rough approximation
$\left\lceil\frac{x-1}{y-x}\right\rceil x\approx\frac{b/q}{c/q}\frac{b}{q}=\frac{b^{2}}{cq}.$
To minimize this, we want $q$ to be large, and to have $x\geq c+1$, we could take $q\approx\frac{b}{c}$, which would give us $\left\lceil\frac{x-1}{y-x}\right\rceil x\approx b$, so this is promising. However, the floors and ceilings mean that the denominator $(y-x)$ may be closer to $(\frac{c}{q}-2)$, and this $(-2)$ term can really hurt us if $q$ is close to $\frac{c}{2}$. Therefore, when $b$ is close to $\frac{c^{2}}{2}$, we will need to make a clever choice of $q$.
Lemma 3.17.
Fix positive integers $b$ and $c$ with $c\geq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$. There exists a positive integer $q$ such that, letting $x$ and $y$ be as in (3.1), we have $x\geq c+1$, $x<y$, and the inequality (3.3) holds.
Proof.
We first note that
$y-x=\left\lfloor\frac{b+c}{q}\right\rfloor-\left\lceil\frac{b+1}{q}\right\rceil\geq\left(\frac{b+c}{q}-\frac{q-1}{q}\right)-\left(\frac{b+1}{q}+\frac{q-1}{q}\right)=\frac{c}{q}-2+\frac{1}{q}\geq\frac{c}{q}-2$
because the floors and ceilings round by at most $\frac{q-1}{q}$. We also have
$\left\lceil\frac{x-1}{y-x}\right\rceil x=\left\lceil\frac{\left\lceil\frac{b+1}{q}\right\rceil-1}{y-x}\right\rceil\left\lceil\frac{b+1}{q}\right\rceil\leq\left(\frac{b/q}{y-x}+1\right)\left(\frac{b}{q}+1\right).$
Now we will have a few different cases, depending on the value of $b$.
Case 1.
$2c\leq b\leq\frac{c^{2}}{20}$. We let $q=\left\lfloor\frac{b}{c}\right\rfloor$, which means that $\frac{b}{c}-1\leq q\leq\frac{b}{c}$. Then
$x=\left\lceil\frac{b+1}{q}\right\rceil\geq\left\lceil\frac{b+1}{b}c\right\rceil\geq c+1.$
Because $b\geq 2c$, we have $q\geq 2$ and therefore
$q\geq\frac{2}{3}(q+1)\geq\frac{2}{3}\frac{b}{c}.$
Because $b\leq\frac{c^{2}}{20}$, we have $q\leq\frac{b}{c}\leq\frac{c}{20}$, so $y-x\geq\frac{c}{q}-2\geq 18$ and
$y-x\geq\frac{18}{20}(y-x+2)\geq\frac{18}{20}\frac{c}{q}.$
Now using that $c\geq 500$, we have
$\left\lceil\frac{x-1}{y-x}\right\rceil x$ $\leq\left(\frac{b/q}{y-x}+1\right)\left(\frac{b}{q}+1\right)\leq\left(\frac{20}{18}\frac{b}{c}+1\right)\left(\frac{3}{2}c+1\right)\leq(1.12\frac{b}{c}+1)(1.51c)$
$\leq 1.7b+0.51c+c\leq 1.7b+0.255b+c=1.955b+c\leq 2b+c+1.$
Case 2.
$\frac{c^{2}}{20}\leq b\leq\frac{c^{2}}{4.2}$. We let $q=\left\lfloor\frac{b}{c}\right\rfloor$ again, but our estimates will look a bit different. We still have $x\geq c+1$ as before. Because $b\geq\frac{c^{2}}{20}\geq 25c$, we have $q\geq 25$ and therefore
$q=\frac{25}{26}(q+1)\geq\frac{25}{26}\frac{b}{c}.$
Because $b\leq\frac{c^{2}}{4.2}$, we have $q\leq\frac{b}{c}\leq\frac{c}{4.2}$. Now since $y-x$ is an integer and $y-x\geq\frac{c}{q}-2\geq 2.2$, we must have $y-x\geq 3$ and
$y-x\geq\frac{3}{5}(y-x+2)\geq\frac{3}{5}\frac{c}{q}.$
Now using that $c\geq 500$, we have
$\left\lceil\frac{x-1}{y-x}\right\rceil x$ $\leq\left(\frac{b/q}{y-x}+1\right)\left(\frac{b}{q}+1\right)\leq\left(\frac{5}{3}\frac{b}{c}+1\right)\left(\frac{26}{25}c+1\right)\leq(1.67\frac{b}{c}+1)(1.05c)$
$\leq 1.76b+0.05c+c\leq 1.76b+0.002b+c=1.762b+c\leq 2b+c+1.$
Case 3.
$\frac{c^{2}}{4.2}\leq b\leq\frac{c^{2}}{3.4}$. We let $q=\lfloor 0.46\sqrt{b}\rfloor$. Note that $0.487\leq\frac{1}{\sqrt{4.2}}\leq\frac{\sqrt{b}}{c}\leq\frac{1}{\sqrt{3.4}}\leq 0.543$. This means that
$x=\left\lceil\frac{b+1}{q}\right\rceil\geq\frac{b}{q}\geq\frac{b}{0.46\sqrt{b}}>\frac{\sqrt{b}}{0.487}\geq c,$
so $x\geq c+1$ because $x$ is an integer. We also have
$c\geq\sqrt{3.4}\sqrt{b}\geq\frac{\sqrt{3.4}}{0.46}q\geq 4.008q,$
so because $y-x$ is an integer and $y-x\geq\frac{c}{q}-2>2.008$, we must have $y-x\geq 3$ and
$y-x\geq\frac{3}{5}(y-x+2)\geq\frac{3}{5}\frac{c}{q}.$
Now using that $\sqrt{b}\geq 0.487c\geq 0.487\cdot 500\geq 240$, we have $q\geq 0.46\sqrt{b}-1\geq 0.4558\sqrt{b}$ and
$\left\lceil\frac{x-1}{y-x}\right\rceil x$ $\leq\left(\frac{5}{3}\frac{b}{c}+1\right)\left(\frac{b}{q}+1\right)\leq\left(\frac{5}{3\sqrt{3.4}}\sqrt{b}+1\right)\left(\frac{\sqrt{b}}{0.4558}+1\right)$
$\leq(0.904\sqrt{b}+1)(2.194\sqrt{b}+1)\leq(0.9082\sqrt{b})(2.1982\sqrt{b})\leq 1.997b\leq 2b+c+1.$
Case 4.
$\frac{c^{2}}{3.4}\leq b\leq\frac{c^{2}}{2}$. Let $q_{0}=\left\lfloor\frac{\sqrt{b}}{\sqrt{3.5}}\right\rfloor$. We will either take $q=q_{0}$ or $q=q_{0}-1$. We first show that for one of these values of $q$, we will have $y-x\geq 2$. Note that $\frac{1}{\sqrt{3.4}}\leq\frac{\sqrt{b}}{c}\leq\frac{1}{\sqrt{2}}$, so
$c\geq\sqrt{2}\sqrt{b}\geq\sqrt{2}\sqrt{3.5}q_{0}\geq 2.64q_{0}.$
Also note that because $c\geq 500$, we have $\sqrt{b}\geq\frac{500}{\sqrt{3.4}}\geq 270$ and $q_{0}\geq\frac{270}{\sqrt{3.5}}-1\geq 140$. Let us first consider $q=q_{0}$. Use integer division to write $b+1=c_{0}q_{0}+r_{0}$ for some integers $c_{0}$ and $10\leq r_{0}\leq q_{0}-$ 1. If
$r_{0}\geq 0.38q_{0}\geq 0.37q_{0}+1,$
this will mean that $c+r_{0}-1\geq 2.64q_{0}+0.37q_{0}=3.01q_{0}>3q_{0}$. Therefore,
$b+c=b+1+c-1=c_{0}q_{0}+(c+r_{0}-1)>(c_{0}+3)q_{0},$
so at least three multiples of $q_{0}$, namely $(c_{0}+1)q_{0}$, $(c_{0}+2)q_{0}$, and $(c_{0}+3)q_{0}$, lie strictly between $b+1$ and $b+c$. This means that $x\leq c_{0}+1$ and $y\geq c_{0}+3$, so $y-x\geq 2$.
Now if $r_{0}<0.38q_{0}$, we will consider $q=q_{0}-1$. We have $\sqrt{3.5}q_{0}\leq\sqrt{b}\leq\sqrt{3.5}(q_{0}+1)$, so
$3.53.5q_{0}^{2}\leq b\leq 3.5q_{0}^{2}+7q_{0}+$ 3.5 .
Now the integer $c_{0}$ satisfies $b+1-q_{0}\leq c_{0}q_{0}\leq b+1$, so using that $q_{0}\geq 140$, we have
$3.49q_{0}^{2}\leq 3.5q_{0}^{2}-q_{0}\leq b+1-q_{0}\leq c_{0}q_{0}\leq b+1\leq 3.5q_{0}^{2}+7q_{0}+4.5\leq 3.56q_{0}^{2}.$
This means that in the integer division of $c_{0}$ by $q_{0}$, we have
$c_{0}=3q_{0}+r_{1},\text{ where }0.49q_{0}\leq r_{1}\leq 0.56q_{0}.$
This allows us to calculate the integer division of $b+1$ by $q=q_{0}-1$ as
$b+1=c_{0}q_{0}+r_{0}=c_{0}(q_{0}-1)+3q_{0}+r_{0}+r_{1}=(c_{0}+3)(q_{0}-1)+(r_{0}+r_{1}+3),$
where the remainder $r=r_{0}+r_{1}+3$ satisfies
$0.49(q_{0}-1)\leq r_{1}\leq r\leq 0.38q_{0}+0.56q_{0}+3=0.94q_{0}+3\leq 0.97(q_{0}-1).$
This means that $c+r-1\geq 2.64q_{0}+0.49(q_{0}-1)-1>3(q_{0}-1)$. Therefore,
$b+c=b+1+c-1=(c_{0}+3)(q_{0}-1)+c+r-1>(c_{0}+6)(q_{0}-1),$
so at least three multiples of $(q_{0}-1)$, namely $(c_{0}+4)(q_{0}-1)$, $(c_{0}+5)(q_{0}-1)$, and $(c_{0}+6)(q_{0}-1)$, lie strictly between $(b+1)$ and $(b+c)$. This means that $x\leq c_{0}+4$ and $y\geq c_{0}+6$, so $y-x\geq 2$. This proves our claim that one of $q=q_{0}$ or $q=q_{0}-1$ will have $y-x\geq 2$.
Now taking either $q=q_{0}$ or $q=q_{0}-1$ so that $y-x\geq 2$, we have
$x=\left\lceil\frac{b+1}{q}\right\rceil\geq\frac{b}{q_{0}}\geq\sqrt{3.5}\sqrt{b}>\sqrt{3.4}\sqrt{b}\geq c,$
so $x\geq c+1$ because $x$ is an integer. Because $\sqrt{b}\geq 270$, we have
$q\geq q_{0}-1\geq\frac{\sqrt{b}}{\sqrt{3.5}}-2\geq\frac{\sqrt{b}}{\sqrt{3.8}}.$
Finally, we have
$\left\lceil\frac{x-1}{y-x}\right\rceil x$ $\leq\left(\frac{b}{2q}+1\right)\left(\frac{b}{q}+1\right)\leq\left(\frac{\sqrt{3.8}b}{2\sqrt{b}}+1\right)\left(\frac{\sqrt{3.8}b}{\sqrt{b}}+1\right)$
$\leq(0.975\sqrt{b}+1)(1.95\sqrt{b}+1)\leq(0.979\sqrt{b})(1.954\sqrt{b})\leq 1.92b\leq 2b+c+1.$
This completes the proof. ∎
Now we can prove the remaining Part 3 of Theorem 3.1.
Proposition 3.18.
If $c\geq 2$ and $2c\leq b\leq\frac{c^{2}}{2}$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
This follows directly from Proposition 3.13, Proposition 3.15, Lemma 3.14, and Lemma 3.17. ∎
We also prove that spiders with four legs are not $e$-positive.
Proof of Theorem 1.6. .
Because $S$ has four legs, we have $c\geq c_{1}+1\geq 2$. If $b\leq\frac{c^{2}}{2}$, the result follows from Parts 1, 2, and 3 of Theorem 3.1. If $b\geq\frac{c^{2}}{2}$, then
$n\geq 2b+c+1\geq c^{2}+c+1,$
and Zheng showed [spiderepos, Corollary 4.6] that such a spider $S$ is not $e$-positive. ∎
In order to prove Conjecture 1.1, it now remains to consider trees with a vertex of degree $4$ that is adjacent to a leaf, such as the spider $S(6,4,1,1)$. It turns out that $S(6,4,1,1)$ has a connected partition of every type $\lambda\vdash 13$, which shows that Wolfgang’s necessary condition will not be enough. We conclude by showing that there are infinitely many such examples.
Figure 3. The spider $S(6,4,1,1)$ has a connected partition of every type $\lambda\vdash 13$
$S(6,4,1,1)$
Proposition 3.19.
Let $m\geq 1$ be an integer. The spider $S(6m,6m-2,1,1)$ has a connected partition of type $\lambda$ for every $\lambda\vdash 12m+1$.
Proof.
We claim that if $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$ has a rearrangement $\alpha$ with $6m-1,6m\notin\text{set}(\alpha)$, then $S(6m,6m-2,1,1)$ has a connected partition of type $\lambda$. Label the vertices on the leg of length $26m-$ 2 as $u_{1},\ldots,u_{6m-2}$, moving in from the leaf, and label the vertices on the leg of length $6m$ as $v_{1},\ldots,v_{6m}$ similarly. Suppose that $\alpha$ is a rearrangement of $\lambda$ that satisfies
$\alpha_{1}+\cdots+\alpha_{i}\leq 6m-2\text{ and }\alpha_{1}+\cdots+\alpha_{i+1}\geq 6m+1.$
Then we can partition the vertices $\{u_{1},u_{2},\ldots,u_{\alpha_{1}+\cdots+\alpha_{i}}\}$ into paths of sizes $\alpha_{1},\ldots,\alpha_{i}$, we can partition the vertices $\{v_{1},v_{2},\ldots,v_{\alpha_{i+2}+\cdots+\alpha_{\ell}}\}$ into paths of sizes $\alpha_{i+2},\ldots,\alpha_{\ell}$, and the remaining vertices form a connected set of size $\alpha_{i+1}$. This establishes our claim.
Note that if $\lambda$ has at least two $1$’s or if $\lambda=(2,2,\ldots,2,1)$, then $S(6m,6m-2,1,1)$ has a connected partition of type $\lambda$, so we may assume that $\lambda$ has at most one $1$ and has a part $x\geq 3$. By our claim, it suffices to find a rearrangement $\alpha$ of $\lambda$ with $6m-1,6m\notin\text{set}(\alpha)$. Fix a rearrangement $\beta$ of $\lambda$. We may assume that $16m-$ 1 or $6m\in\text{set}(\beta)$, otherwise we could take $\alpha=\beta$. Letting $\beta^{*}=(\beta_{\ell},\ldots,\beta_{1})$ denote the reverse composition, we may assume that $16m-$ 1 or $6m\in\text{set}(\beta^{*})$, otherwise we could take $\alpha=\beta^{*}$. Now there are only a few cases to consider.
If $6m\in\text{set}(\beta)$ and $6m\in\text{set}(\beta^{*})$, we must have $\beta_{1}+\cdots+\beta_{i}=6m$, $\beta_{i+1}=1$, and $\beta_{i+2}+\cdots+\beta_{\ell}=6m$ for some $i$. By reversing, we may assume without loss of generality that some part $\beta_{k}=x\geq 3$ for $1\leq k\leq i$. Then the composition
$\alpha=(\beta_{1},\ldots,\beta_{k-1},\beta_{k+1},\ldots,\beta_{i},1,x,\beta_{i+2},\ldots,\beta_{\ell})$
has $\text{set}(\alpha)=\{\ldots,6m-x,6m-x+1,6m+1,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$.
If $6m-1\in\text{set}(\beta)$, $6m\notin\text{set}(\beta)$, and $6m\in\text{set}(\beta^{*})$, we must have $\beta_{1}+\cdots+\beta_{i}=6m-1$, $\beta_{i+1}=2$, and $\beta_{i+2}+\cdots+\beta_{\ell}=6m$ for some $i$. If there is some $\beta_{k}=x\geq 3$ for $1\leq k\leq i$, then the composition
$\alpha=(\beta_{1},\ldots,\beta_{k-1},\beta_{k+1},\ldots,\beta_{i},2,x,\beta_{i+2},\ldots,\beta_{\ell})$
has $\text{set}(\alpha)=\{\ldots,6m-x-1,6m-x+1,6m+1,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$. Otherwise, the parts $(\beta_{1},\ldots,\beta_{i})$ consist of a single $1$ and $(3m-1)$ $2$’s, some $\beta_{k}=x\geq 3$ for $i+2\leq k\leq\ell$, and the composition
$\alpha=(2,2,\ldots,2,2,x,1,2,\beta_{i+2},\ldots,\beta_{k-1},\beta_{k+1},\ldots,\beta_{\ell})$
has $\text{set}(\alpha)=\{\ldots,6m-2,6m-2+x,6m-1+x,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$. If $6m\in\text{set}(\beta)$, $6m-1\in\text{set}(\beta^{*})$, and $6m\notin\text{set}(\beta^{*})$, we can reverse $\beta$ to reduce to the previous case.
Finally, if $6m-1\in\text{set}(\beta)$, $6m\notin\text{set}(\beta)$, $6m-1\in\text{set}(\beta^{*})$, and $6m\notin\text{set}(\beta^{*})$, then we must have $\beta_{1}+\cdots+\beta_{i}=6m-1$, $\beta_{i+1}=3$, and $\beta_{i+2}+\cdots+\beta_{\ell}=6m-1$ for some $i$. If $\beta_{k}=1$ for some $1\leq k\leq i$, then the composition
$\alpha=(\beta_{1},\ldots,\beta_{k-1},\beta_{k+1},\ldots,\beta_{i},3,1,\beta_{i+2},\ldots,\beta_{\ell})$
has $\text{set}(\alpha)=\{\ldots,6m-2,6m+1,6m+2,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$. If $\beta_{k}=y\geq 4$ for some $1\leq k\leq i$, then the composition
$\alpha=(\beta_{1},\ldots,\beta_{k-1},\beta_{k+1},\ldots,\beta_{i},3,y,\beta_{i+2},\ldots,\beta_{\ell})$
has $\text{set}(\alpha)=\{\ldots,6m-1-y,6m+2-y,6m+2,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$. Now we may assume that every $1\leq k\leq i$ has $\beta_{k}=2$ or $3$, and by reversing, that every $i+2\leq k\leq\ell$ also has $\beta_{k}=2$ or $3$. Because $16m-$ 1 is not divisible by either $2$ or $3$, there must be some $1\leq k\leq i$ with $\beta_{k}=3$ and some $i+2\leq k^{\prime}\leq\ell$ with $\beta_{k^{\prime}}=2$. Now the composition
$\alpha=(\beta_{1},\ldots,\beta_{k-1},\beta_{k+1},\ldots,\beta_{i},2,3,3,\beta_{i+2},\ldots,\beta_{k^{\prime}-1},\beta_{k^{\prime}+1},\ldots,\beta_{\ell})$
has $\text{set}(\alpha)=\{\ldots,6m-4,6m-2,6m+1,6m+4,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$. This completes the proof. ∎
4. Acknowledgments
The author would like to think José Aliste-Prieto, Richard Stanley, Stephanie van Willigenburg, and Kai Zheng for helpful discussions.
[…]

### v2

@@
-TreeFive.bib
+Graphs missing a connected partition
 Abstract.
-We prove that a tree with a vertex of degree at least five must be missing a connected partition of some type and therefore its chromatic symmetric function cannot be $e$-positive.
-We prove that this also holds for a tree with a vertex of degree four as long as it is not adjacent to any leaf.
+We prove that a graph with a cut vertex whose deletion produces at least five connected components must be missing a connected partition of some type.
+We prove that this also holds if there are four connected components that each have at least two vertices.
+In particular, the chromatic symmetric function of such a graph cannot be $e$-positive.
 This brings us very close to the conjecture by Dahlberg, She, and van Willigenburg of non-$e$-positivity for all trees with a vertex of degree at least four.
@@
 Key words and phrases:
-chromatic symmetric function, connected partition, cut vertex, e-positive, proper colouring, spider graph, Stanley–Stembridge conjecture, tree isomorphism conjecture
+chromatic symmetric function, connected partition, cut vertex, e-positive, graph partition, Stanley–Stembridge conjecture, tree isomorphism conjecture
 2020 Mathematics Subject Classification:
-Primary 05E05; Secondary 05C05, 05C15
+Primary 05C70; Secondary 05E05
 Department of Mathematics, MIT
 1. Introduction
-When Stanley [chromsym] defined the chromatic symmetric function $X_{G}$ of a graph $G$, he posed two monumental problems that have garnered tremendous interest.
-The first is the Stanley–Stembridge conjecture [stanstem], which asserts that $X_{G}$ is $e$-positive whenever $G$ is the incomparability graph of a $(\bm{3}+\bm{1})$-free poset.
-Many authors proved $e$-positivity for particular subclasses of such graphs [csfvertex, chrombounce3, chromstoe, chromqrook, trilad, chromnsym, qforesttriples] and studied connections to Hessenberg varieties [chromhesssplitting, dothessenberg, hessenberghopf, stanstemhess, chromquasihessenberg] and Hecke algebras [chromhecke, chromcharacters].
-The second is the tree isomorphism conjecture, which asserts that trees $T_{1}$ and $T_{2}$ can have $X_{T_{1}}=X_{T_{2}}$ only if they are isomorphic.
-This is intriguing because all trees with $n$ vertices have the same chromatic polynomial.
-Many authors proved that the chromatic symmetric function determines certain invariants and therefore distinguishes particular subclasses of trees [propcatchrom, chromtreepolyinv, markedgraphs, propqcatchrom, algochromtrees, catchromiso, distinguishingtrees, chromequal].
-Dahlberg, She, and van Willigenburg [chromspostrees] considered a hybrid of these two questions by asking which trees $T$ have that $X_{T}$ is $e$-positive.
-Many authors [eposclawcon, spider4m22m1, spiderbroom, spiderepos] proved $e$-positivity or non-$e$-positivity for particular subclasses of trees.
+Graph partitioning is a fundamentally important problem in computer science.
+For example, the max-flow min-cut theorem [7] is a cornerstone result in optimization theory.
+There are numerous practical applications for finding communities in large networks and graph partitions can allow for parallelization of algorithms to reduce computational complexity.
+Surveys of graph partitioning problems and algorithms can be found in [1, 3, 11, 12].
+Let $G=(V,E)$ be an $n$-vertex graph and let $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$ be an integer partition of $n$.
+A connected partition of $G$ of type $\lambda$ is a partition of the vertices $V$ into connected subsets of sizes $\lambda_{1},\ldots,\lambda_{\ell}$.
+A more precise definition and examples will be given in Section 2.
+Many authors studied problems of finding connected partitions that satisfy certain additional constraints or optimizations [2, 4, 8, 9, 10, 13].
+In this paper, we consider the problem of finding an integer partition $\lambda$ for which $G$ is missing a connected partition of type $\lambda$.
+One application of this is to non-$e$-positivity of chromatic symmetric functions.
+Stanley [14] defined the chromatic symmetric function of a graph $G=(V,E)$ to be
+$X_{G}=\sum_{\begin{subarray}{c}\kappa:V\to\{1,2,3,\ldots\}\\ \text{ if }ij\in E,\text{ then }\kappa(i)\neq\kappa(j)\end{subarray}}\prod_{v\in V}x_{\kappa(v)}$
+and we say that $X_{G}$ is $e$-positive if it can be written as a nonnegative linear combination in the basis of elementary symmetric functions, defined by
+$e_{\lambda}=e_{\lambda_{1}}\cdots e_{\lambda_{\ell}},\text{ where }e_{k}=\sum_{i_{1}<\cdots<i_{k}}x_{i_{1}}\cdots x_{i_{k}}.$
+It is also common to say that $G$ itself is $e$-positive. Wolfgang found a powerful necessary condition for $e$-positivity.
+Lemma 1.1.
+[18, Proposition 1.3.3] If a connected graph $G$ is missing a connected partition of some type $\lambda\vdash n$, then it is not $e$-positive.
+Dahlberg, She, and van Willigenburg [6] were interested in determining which trees are $e$-positive.
+Many authors [5, 6, 16, 17, 19] proved $e$-positivity or non-$e$-positivity for particular subclasses of trees.
 Dahlberg, She, and van Willigenburg checked the following conjecture for all trees with at most $12$ vertices.
 Conjecture 1.2.
-[chromspostrees, Conjecture 6.1] If $T$ is a tree with a vertex of degree $d\geq 4$, then $X_{T}$ is not $e$-positive.
-A powerful tool for studying $e$-positivity is the following necessary condition by Wolfgang.
-Proposition 1.2.
-[conparts, Proposition 1.3.3] A connected graph $G$ must have a connected partition of every type in order for $X_{G}$ to be $e$-positive.
-Note that this condition is not sufficient because for example, for every $m\geq 1$ the spider graph $G=S(2m+2,1,1)$ has a connected partition of every type but $X_{G}$ is not $e$-positive.
-Dahlberg, She, and van Willigenburg proved the following result.
+[6, Conjecture 6.1] If $T$ is a tree with a vertex of degree at least $4$, then $T$ is not $e$-positive.
+They proved the following result.
 Theorem 1.3.
-[chromspostrees, Theorem 4.1] If $G$ is an $n$-vertex connected graph with a cut vertex whose deletion produces a graph with $d\geq\log_{2}n+1$ connected components, then $G$ is missing a connected partition of some type and therefore $X_{G}$ is not $e$-positive.
-In particular, if $T$ is an $n$-vertex tree with a vertex of degree $d\geq\log_{2}n+1$, then $X_{T}$ is not $e$-positive. Zheng strengthened this result considerably.
+[6, Theorem 4.1] If $G$ is an $n$-vertex graph with a cut vertex whose deletion produces $d\geq\log_{2}n+1$ connected components, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
+In particular, if $T$ is an $n$-vertex tree with a vertex of degree $d\geq\log_{2}n+1$, then $T$ is not $e$-positive. Zheng strengthened this result considerably.
 Theorem 1.4.
-[spiderepos, Theorem 3.13] If $G$ is a connected graph with a cut vertex whose deletion produces a graph with at least $6$ connected components, then $G$ is missing a connected partition of some type.
+[19, Theorem 3.13] If $G$ is a graph with a cut vertex whose deletion produces at least $6$ connected components, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
 Our main result lowers this bound further.
 Theorem 1.5.
-Let $G$ be a connected graph with a cut vertex whose deletion produces a graph with either at least $5$ connected components, or $4$ connected components that each have at least two vertices.
-Then $G$ is missing a connected partition of some type.
-In particular, if $T$ is a tree with a vertex of degree at least $5$, or with a vertex of degree $4$ that is not adjacent to any leaf, then $T$ is missing a connected partition of some type and therefore $X_{T}$ is not $e$-positive.
-Note that there exists a tree $T$ with a vertex of degree $4$ that is adjacent to a leaf, and which has a connected partition of every type.
+Let $G$ be a graph with a cut vertex whose deletion produces either at least $5$ connected components, or $4$ connected components that each have at least two vertices.
+Then $G$ is missing a connected partition of some type $\lambda\vdash n$.
+In particular, if $T$ is a tree with a vertex of degree at least $5$, or with a vertex of degree $4$ that is not adjacent to any leaf, then $T$ is not $e$-positive.
+We will show in Proposition 3.18 that there exists an infinite family of trees with a vertex of degree $4$ that is adjacent to a leaf, and which have a connected partition of every type $\lambda\vdash n$.
 Therefore, in this sense, Theorem 1.5 cannot be improved.
-In fact, we will show in Proposition 3.19 that for every $m\geq 1$ the spider graph $S(6m,6m-2,1,1)$ has a connected partition of every type, so there are infinitely many such examples.
 This means that to prove non-$e$-positivity of trees with a vertex of degree $4$ in general, we need to know more about the tree to calculate particular coefficients.
-In the specific case of spiders, we will prove that Conjecture 1.1 holds.
+In the specific case of spiders, which are graphs $S(\mu)$ formed by joining paths of lengths $\mu_{1},\ldots,\mu_{\ell}$, the legs of the spider, at a single vertex, we will prove that Conjecture 1.2 holds.
 Theorem 1.6.
-If $S$ is a spider with four legs, then $X_{S}$ is not $e$-positive.
-Note that there exist infinite families of spiders with three legs, such as $S(m+1,m,1)$ for $m\geq 1$, that have $e$-positive chromatic symmetric function [chromspostrees, Theorem 3.2].
-In the next section, we define chromatic symmetric functions, connected partitions, and spider graphs. In Section 3, we prove Theorem 1.5 by analyzing partial sums of rearrangements of partitions.
+Spiders with four legs are not $e$-positive.
 2. Background
-A composition is a finite sequence of positive integers $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$, which we call the parts of $\alpha$. We denote by
-$\text{set}(\alpha)=\{\alpha_{1},\alpha_{1}+\alpha_{2},\ldots,\alpha_{1}+\cdots+\alpha_{\ell-2},\alpha_{1}+\cdots+\alpha_{\ell-1}\}$
-the set of (proper) partial sums of $\alpha$.
-A partition is a composition $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$ that also satisfies $\lambda_{1}\geq\cdots\geq\lambda_{\ell}$.
-We write $\lambda\vdash n$ if $\lambda_{1}+\cdots+\lambda_{\ell}=n$.
-We say that $\alpha$ is a rearrangement of $\lambda$ if every part occurs the same number of times in $\alpha$ and in $\lambda$.
-Example 2.1.
-The composition $\alpha=(3,4,2,4)$ has $\text{set}(\alpha)=\{3,7,9\}$ and is a rearrangement of the partition $\lambda=(4,4,3,2)\vdash 13$.
-Let $G=(V,E)$ be a graph with $n$ vertices.
-A proper colouring of $G$ is a function $\kappa:V\to\{1,2,3,\ldots\}$ such that whenever vertices $u,v\in V$ are adjacent, we must have $\kappa(u)\neq\kappa(v)$.
-The chromatic symmetric function of $G$, defined by Stanley [chromsym, Definition 2.1] is the formal power series in infinitely many commuting variables $(x_{1},x_{2},x_{3},\ldots)$ given by
-$X_{G}=\sum_{\kappa\text{ proper colouring }}\prod_{v\in V}x_{\kappa(v)}.$
-We can write $X_{G}$ uniquely as a linear combination of elementary symmetric functions $e_{\lambda}$, which are indexed by partitions $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})\vdash n$ and defined by
-$e_{\lambda}=e_{\lambda_{1}}\cdots e_{\lambda_{\ell}},\text{ where }e_{k}=\sum_{i_{1}\leq\cdots\leq i_{k}}x_{i_{1}}\cdots x_{i_{k}}.$
-We say that $X_{G}$ is $e$-positive if all of the coefficients $c_{\lambda}$ in the expansion $X_{G}=\sum_{\lambda\vdash n}c_{\lambda}e_{\lambda}$ are nonnegative.
-It is also common to say that $G$ itself is $e$-positive.
-A connected partition of $G$ of type $\lambda\vdash n$ is a collection $\mathcal{S}=\{S_{1},\ldots,S_{\ell}\}$ of pairwise disjoint subsets of $V$ such that
-  • $\circ$
-    $\bigcup_{i=1}^{\ell}S_{i}=V$,
+[…]
+$S(6,4,1,1)=$
+A graph $G=(V,E)$ consists of a set $V$ of vertices and a set $E$ of edges, which are sets of two vertices.
+In particular, we do not allow loops, which are edges joining a vertex to itself, or parallel edges, which are edges joining the same pair of vertices.
+We will always use $n$ to denote the order of $G$, which is the number of vertices.
+For a subset $S\subseteq V$, the induced subgraph is $G[S]=(S,\{e=vw\in E:\ v,w\in S\})$.
+We say that $G$ is connected if for every $v,w\in V$, there is a path from $v$ to $w$, which is a sequence of vertices $(v=v_{1},v_{2}\ldots,v_{k-1},v_{k}=w)$ with every $v_{i}v_{i+1}\in E$.
+A connected component of $G$ is a maximal subset $S\subseteq V$ such that $G[S]$ is connected.
+A vertex $v\in V$ of a connected graph is a cut vertex if the graph $G\setminus v$ obtained by deleting $v$ and all incident edges is not connected.
+A tree $T=(V,E)$ is a minimally connected graph, meaning that for every $e\in E$, the graph $(V,E\setminus\{e\})$ is not connected.
+The degree of a vertex $v$ is the number of edges incident to it and a leaf of a tree is a vertex of degree one.
+Note that in a tree, every vertex is a leaf or a cut vertex.
+If $v$ is a cut vertex of degree $d$, then $T\setminus v$ has $d$ connected components.
+A composition of $n$ is a sequence of positive integers $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$ with sum $n$.
+An (integer) partition of $n$ is a weakly decreasing composition $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$ of $n$.
+These integers $\lambda_{i}$ are called the parts of $\lambda$.
+We write $\lambda\vdash n$ to mean that $\lambda$ is an integer partition of $n$.
+We now define the main object of study in this paper.
+Definition 2.1.
+Let $G=(V,E)$ be an $n$-vertex graph and let $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})\vdash n$.
+A connected partition of $G$ of type $\lambda$ is a set partition $\mathcal{S}=\{S_{1},\ldots,S_{\ell}\}$ of $V$ so that
   • $\circ$
@@
     $|S_{i}|=\lambda_{i}$ for every $1\leq i\leq\ell$.
-Proposition 2.2.
-[conparts, Proposition 1.3.3] Let $G$ be a connected $e$-positive graph with $n$ vertices. Then for every partition $\lambda\vdash n$, there exists a connected partition of $G$ of type $\lambda$.
-A tree is a connected acyclic graph.
-A leaf of a tree is a vertex of degree one.
-A cut vertex of a connected graph $G$ is a vertex whose deletion results in a disconnected graph.
-For a partition $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$, we denote by $S(\lambda)$ the graph consisting of paths of lengths $\lambda_{1},\ldots,\lambda_{\ell}$, called legs, joined at a common vertex.
-Graphs of the form $S(\lambda)$ are called spiders.
+We say that $G$ is missing a connected partition of type $\lambda$ if no such $\mathcal{S}$ exists.
 Example 2.2.
-Figure 1 shows some spiders and their chromatic symmetric functions.
-The spider $S(1,1,1)$ is missing a connected partition of type $(2,2)$ and is indeed not $e$-positive.
-The spider $S(3,2,1)$ has a connected partition of type $\lambda$ for every $\lambda\vdash 7$ and is $e$-positive.
-The spider $S(4,1,1)$ has a connected partition of type $\lambda$ for every $\lambda\vdash 7$ but is not $e$-positive, showing that Wolfgang’s necessary condition for $e$-positivity is not sufficient.
-Figure 1. Some spiders and their chromatic symmetric functions
-$S(1,1,1)$         $S(3,2,1)$         $S(4,1,1)$
-$X_{S(1,1,1)}$ $=e_{211}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-2e_{22}}+5e_{31}+4e_{4}$
-$X_{S(3,2,1)}$ $=e_{2221}+2e_{3211}+5e_{322}+4e_{331}+12e_{421}+5e_{43}+4e_{511}+13e_{52}+11e_{61}+7e_{7}$
-$X_{S(4,1,1)}$ $=e_{2221}+4e_{3211}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-3e_{322}}+10e_{331}+10e_{421}+17e_{43}+4e_{511}+3e_{52}+11e_{61}+7e_{7}$
+The spider graph $G=S(6,4,1,1)$ is shown in Figure 1.
+We have indicated a connected partition of $G$ of type $\lambda=(5,4,3,1)\vdash 13$.
+We will see in Proposition 3.18 that $G$ has a connected partition of every type $\lambda\vdash 13$.
+But by Theorem 1.6, $G$ still fails to be $e$-positive.
+This shows that Wolfgang’s necessary condition for $e$-positivity is not sufficient.
+Example 2.3.
+If $n$ is even, a connected partition of $G=(V,E)$ of type $\lambda=(2,2,\ldots,2,2)$ is equivalent to a perfect matching of $G$, which is a subset $U\subseteq E$ such that every vertex is in exactly one edge of $U$.
 3. Proof of main result
+Figure 2. A graph $G$ with a cut vertex $v$ joining at least $3$ connected components
+[…]
 We first fix notation that we will use throughout this section.
-Let $G$ be an $n$-vertex connected graph with a cut vertex $v$ whose deletion produces a graph with at least $3$ connected components, and denote their sizes
+Notation 3.1.
+Let $G$ be an $n$-vertex connected graph with a cut vertex $v$ whose deletion produces at least $3$ connected components, and denote their orders
 $a\geq b\geq c_{1}\geq\cdots\geq c_{k}\geq 1,\text{ where }k\geq 1.$
@@
 $n=a+b+c+1\geq 2b+c+1.$
-We now state our main result using this notation.
+We will always use $\lambda$ to denote an integer partition of $n$.
+We now restate our main result using this notation.
 Theorem 3.2.
-Suppose that $c\geq 2$. If any of the following hold, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
+Let $G$, $a$, $b$, $c$, $c_{1}$, and $n$ be as in Notation 3.1. Suppose that $c\geq 2$. If any of the following hold, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
   1. (1)
@@
   2. (2)
-    We have $c\geq c_{1}+1$ and $b=2c-1$.
+    We have b=2c1b=2c-1 and $c\geq c_{1}+1$.
   3. (3)
@@
   4. (4)
-    We have $c\geq c_{1}+2$ and $b\geq\frac{c^{2}}{2}$.
-In particular, if $c\geq c_{1}+2$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
+    We have $b\geq\frac{c^{2}}{2}$ and $c\geq c_{1}+2$.
+In particular, if $c\geq c_{1}+2$, then no matter which range $b$ lies in, we can conclude that $G$ is missing a connected partition of some type $\lambda\vdash n$.
+The proof of Theorem 3.2 is structured as follows.
+We begin by reformulating connected partitions of type $\lambda$ in terms of partial sums of rearrangements of $\lambda$ (Lemma 3.7), which we use to prove Part (4).
+We then describe a strategy for choosing the parts of $\lambda$ (Lemma 3.9), which will prove Parts (1) and (2).
+To prove Part (3), we will need a technical result to establish an inequality (Lemma 3.17).
+Our estimates will only be valid if $c$ is sufficiently large, namely $c\geq 500$, but we can check the cases with $2\leq c\leq 499$ by computer (Lemma 3.15).
+Sage code used to check these cases is provided in the Appendix.
 Remark 3.3.
 The condition that $c\geq c_{1}+2$ is precisely that the deletion of $v$ produces either at least $5$ connected components, or $4$ connected components that each have at least two vertices, so Theorem 3.2 implies Theorem 1.5.
-We cannot remove this hypothesis because the spider graph $S(6,4,1,1)$ has a connected partition of every type $\lambda\vdash 13$.
+We cannot remove this hypothesis because the spider graph $S(6,4,1,1)$ from Figure 1 has a connected partition of every type $\lambda\vdash 13$.
+Incidentally, it will follow from Theorem 1.6 that the graph $S(6,4,1,1)$ is not $e$-positive.
 Remark 3.4.
@@
 Remark 3.5.
-We cannot remove the hypothesis that $c\geq 2$ because the family of spider graphs $S(m+1,m,1)$ all have a connected partition of every type $\lambda\vdash 2m+3$ and are in fact $e$-positive.
+We cannot remove the hypothesis that $c\geq 2$ because the family of spider graphs $S(m+1,m,1)$ all have a connected partition of every type $\lambda\vdash 2m+3$ and are in fact $e$-positive [5, Theorem 3.2].
 We begin by describing another way to think about connected partitions.
+A composition $\alpha=(\alpha_{1},\ldots,\alpha_{\ell})$ is a rearrangement of $\lambda$ if each positive integer occurs the same number of times in $\alpha$ and in $\lambda$.
+We denote by
+$\text{set}(\alpha)=\{\alpha_{1},\alpha_{1}+\alpha_{2},\ldots,\alpha_{1}+\cdots+\alpha_{\ell-2},\alpha_{1}+\cdots+\alpha_{\ell-1}\}$
+the set of (proper) partial sums of $\alpha$.
+Example 3.6.
+The composition $\alpha=(3,1,5,4)$ is a rearrangement of $\lambda=(5,4,3,1)$ and has set of partial sums $\text{set}(\alpha)=\{3,4,9\}$.
 Lemma 3.7.
-Let $\lambda\vdash n$ be a partition with all parts at least $(c_{1}+1)$.
-If every rearrangement $\alpha$ of $\lambda$ has a partial sum $s\in\text{set}(\alpha)$ that lies in the interval $I=\{b+1,b+2,\ldots,b+c-1,b+c\}$, then $G$ has no connected partition of type $\lambda$.
+Let $G$, $a$, $b$, $c$, $c_{1}$, $n$, and $v$ be as in Notation 3.1.
+Let $\lambda$ be a partition of $n$ with all parts at least $(c_{1}+1)$.
+If every rearrangement $\alpha$ of $\lambda$ has a partial sum $s\in\text{set}(\alpha)$ that lies in the interval $I=\{b+1,\ldots,b+c\}$, then $G$ is missing a connected partition of type $\lambda$.
 Proof.
@@
 We will find a rearrangement $\alpha$ of $\lambda$ that has no partial sum in $I$.
-Let $A,B,C_{1},\ldots,C_{k}$ be the connected components of $G\setminus v$ whose sizes are respectively $a\geq b\geq c_{1}\geq\cdots\geq c_{k}$ and let $S\in\mathcal{S}$ be the subset with $v\in S$.
+Let $A,B,C_{1},\ldots,C_{k}$ be the connected components of $G\setminus v$ whose orders are respectively $a\geq b\geq c_{1}\geq\cdots\geq c_{k}$ and let $S\in\mathcal{S}$ be the subset with $v\in S$.
 If any $C_{i}\nsubseteq S$, then letting $u\in C_{i}\setminus S$, the subset $S^{\prime}\in\mathcal{S}$ with $u\in S^{\prime}$ would have $|S^{\prime}|\leq c_{i}\leq c_{1}$, which is impossible because every part of $\lambda$ is at least $(c_{1}+1)$.
@@
 Therefore, no partial sum of $\alpha$ lies in the interval $I=\{b+1,\ldots,b+c\}$. ∎
-Figure 2. A graph $G$ with a cut vertex $v$ joining at least $3$ connected components
-[…]
-We can now prove Part 4 of Theorem 3.1, the case where $b\geq\frac{c^{2}}{2}$.
-Proposition 3.6.
-If $c\geq c_{1}+2$ and $b\geq\frac{c^{2}}{2}$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
-Proof.
-The idea is to take a partition $\lambda\vdash n$ with all parts equal to $(c-1)$ or $c$.
-The well-known Frobenius coin problem tells us that every integer $n\geq(c-1)(c-2)$ can be written in the form $n=a_{1}c+a_{2}(c-1)$ for some integers $a_{1},a_{2}\geq 0$.
-Because $b\geq\frac{c^{2}}{2}$, we have
+We also use the following result, the two-coin solution to the Frobenius coin problem, which gives a sufficient condition to find a partition $\lambda$ of $n$ with specified parts.
+Theorem 3.8.
+[15] Let $x$ and $y$ be positive integers with $\gcd(x,y)=1$. Then every integer $n\geq(x-1)(y-1)$ can be written in the form $n=a_{1}x+a_{2}y$ for some integers $a_{1},a_{2}\geq 0$.
+We can now prove Part (4) of Theorem 3.2, the case where $c\geq c_{1}+2$ and $b\geq\frac{c^{2}}{2}$.
+Proof of Part (4) of Theorem 3.2. .
+The idea is to take a partition $\lambda$ of $n$ with parts $c$ or (c1)(c-1). Because $b\geq\frac{c^{2}}{2}$, we have
 $n\geq 2b+c+1\geq c^{2}+c+1\geq(c-1)(c-2),$
-so we can write $n=a_{1}c+a_{2}(c-1)$ and take $\lambda$ to be $a_{1}$ $c$’s followed by $a_{2}$ $(c-1)$’s.
+so by Theorem 3.8, we can write $n=a_{1}c+a_{2}(c-1)$ and take $\lambda$ to be $a_{1}$ $c$’s followed by $a_{2}$ (c1)(c-1)’s.
 All parts of $\lambda$ are at least $c-1\geq c_{1}+1$.
 Let $\alpha$ be a rearrangement of $\lambda$.
-Because $c\geq c_{1}+2\geq 3$, we have $\alpha_{1}\leq c\leq\frac{c^{2}}{2}\leq b$, so there exists some maximal $i$ with $\alpha_{1}+\cdots+\alpha_{i}\leq b$, and then
+Because $c\geq c_{1}+2\geq 3$, we have $\alpha_{1}\leq c\leq\frac{c^{2}}{2}\leq b$, so there exists some maximal $i$ with $\alpha_{1}+\cdots+\alpha_{i}\leq b$, and then the partial sum $\alpha_{1}+\cdots+\alpha_{i}+\alpha_{i+1}$ satisfies
 $b+1\leq\alpha_{1}+\cdots+\alpha_{i}+\alpha_{i+1}\leq b+\alpha_{i+1}\leq b+c.$
 Therefore the result follows from Lemma 3.7. ∎
-We now describe another way to ensure that all rearrangements have a partial sum in $I$.
+We now describe another way to ensure that all rearrangements of $\lambda$ have a partial sum in the interval $I=\{b+1,\ldots,b+c\}$.
 Lemma 3.9.
-Let $q$ be a positive integer and let $J=\{x,x+1,\ldots,y-1,y\}$, where
+Let $G$, $a$, $b$, $c$, $c_{1}$, and $n$ be as in Notation 3.1. Let $q$ be a positive integer and let $J=\{x,x+1,\ldots,y-1,y\}$, where
 (3.1) $x=\left\lceil\frac{b+1}{q}\right\rceil\text{ and }y=\left\lfloor\frac{b+c}{q}\right\rfloor.$
-Suppose that $x\geq c_{1}+1$. If $\lambda\vdash n$ is a partition with all parts in the interval $J$, then $G$ has no connected partition of type $\lambda$.
+Suppose that $x\geq c_{1}+1$. If $\lambda$ is a partition of $n$ with all parts in the interval $J$, then $G$ is missing a connected partition of type $\lambda$.
 Proof.
@@
 Lemma 3.10.
-Let $x$ and $y$ be positive integers with $x<y$. Then there exists a partition $\lambda\vdash n$ with all parts in the interval $J=\{x,\ldots,y\}$ as long as
+Let $x$ and $y$ be positive integers with $x<y$. Then there exists a partition $\lambda$ of $n$ with all parts in the interval $J=\{x,\ldots,y\}$ as long as
 (3.2) $n\geq\left\lceil\frac{x-1}{y-x}\right\rceil x.$
@@
 For a positive integer $t$, let $tJ=\{tx,tx+1,\ldots,ty-1,ty\}$.
-First note that every integer $n\in tJ$ has a partition $\lambda\vdash n$ consisting of $t$ integers in $J$.
-Indeed, for $n=ty$ we can take $\lambda=(y,\ldots,y)$, and for $tx\leq n\leq ty-1$ we can write $n=tq+r$ for some integers $x\leq q\leq y-1$ and $10\leq r\leq t-$ 1, and take $\lambda$ to be $r$ $(q+1)$’s followed by $(t-r)$ $q$’s.
+First note that every integer $n\in tJ$ has a partition $\lambda$ consisting of $t$ integers in $J$.
+Indeed, for $n=ty$ we can take $\lambda=(y,\ldots,y)$, and for $tx\leq n\leq ty-1$ we can write $n=tq+r$ for some integers $x\leq q\leq y-1$ and $0\leq r\leq t-1$, and take $\lambda$ to be $r$ $(q+1)$’s followed by (tr)(t-r) $q$’s.
 We now see that as $t$ increases, the intervals $tJ$ become closer to each other.
 Specifically, for $t\geq\left\lceil\frac{x-1}{y-x}\right\rceil\geq\frac{x-1}{y-x}$, we have $ty+1\geq(t+1)x$, which means that as soon as we exit the interval $tJ$, we are already in the interval $(t+1)J$.
-This means that every integer $n$ satisfying (3.2) is in some interval $tJ$ and has a partition $\lambda\vdash n$ with all parts in $J$.
+This means that every integer $n$ satisfying (3.2) is in some interval $tJ$ and has a partition $\lambda$ with all parts in $J$.
 ∎
@@
 We see that once $t\geq\left\lceil\frac{x-1}{y-x}\right\rceil=3$, we no longer have a gap between the intervals $tJ$ and $(t+1)J$, so we are able to express any integer $n\geq 3x=21$.
-Proposition 3.10.
-If $n\geq\left\lceil\frac{b}{c-1}\right\rceil(b+1)$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
+Lemma 3.12.
+Let $G$, $a$, $b$, $c$, $c_{1}$, and $n$ be as in Notation 3.1. If $n\geq\left\lceil\frac{b}{c-1}\right\rceil(b+1)$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
 Proof.
-By Lemma 3.8, there exists a partition $\lambda\vdash n$ with all parts in the interval $J=I=\{b+1,\ldots,b+c\}$.
-All parts of $\lambda$ are at least $b+1\geq c_{1}+1$ and trivially every rearrangement $\alpha$ of $\lambda$ has the partial sum $\alpha_{1}\in\{b+1,\ldots,b+c\}$, so the result follows from Lemma 3.5.
-∎
-We can now prove Parts 1 and 2 of Theorem 3.1.
-Proposition 3.11.
-If $c\geq 2$ and $b\leq 2c-2$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
-Proof.
-If $b\leq 2c-2$, then we have
+By Lemma 3.10, there exists a partition $\lambda$ of $n$ with all parts in the interval $J=\{b+1,\ldots,b+c\}$. Now by Lemma 3.9 with $q=1$, $G$ is missing a connected partition of type $\lambda$. ∎
+We can now prove Part (1) of Theorem 3.2, the case where $b\leq 2c-2$; and Part (2) of Theorem 3.2, the case where $c\geq c_{1}+1$ and b=2c1b=2c-1.
+Proof of Part (1) of Theorem 3.2. .
+Because $b\leq 2c-2$, we have
 $\left\lceil\frac{b}{c-1}\right\rceil(b+1)\leq 2(b+1)\leq 2b+c+1\leq n,$
-so the result follows from Proposition 3.10. ∎
-Proposition 3.12.
-If $c\geq c_{1}+1$ and $b=2c-1$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
-Proof.
-First note that if $c=2$, then $b=3$, $I=\{4,5\}$, and taking $\lambda=(2,\ldots,2)$ if $n$ is even or $\lambda=(3,2,\ldots,2)$ if $n$ is odd, every rearrangement $\alpha$ has $\alpha_{1}+\alpha_{2}=4$ or $5$ and we are done by Lemma 3.5.
+so the result follows from Lemma 3.12. ∎
+Proof of Part (2) of Theorem 3.2. .
+First note that if $c=2$, then $b=3$, and taking $\lambda=(2,\ldots,2)$ if $n$ is even or $\lambda=(3,2,\ldots,2)$ if $n$ is odd, every rearrangement $\alpha$ has the partial sum $\alpha_{1}+\alpha_{2}$ in the interval $I=\{4,5\}$ and we are done by Lemma 3.7.
 Now suppose that $c\geq 3$.
 Let $x=\left\lceil\frac{b+1}{2}\right\rceil=c\geq c_{1}+1$ and $y=\left\lfloor\frac{b+c}{2}\right\rfloor=\left\lfloor\frac{3c-1}{2}\right\rfloor$.
-By Lemma 3.8, we can find a partition $\lambda\vdash n$ with all parts in $J=\{x,\ldots,y\}$ because
+By Lemma 3.10, we can find a partition $\lambda$ of $n$ with all parts in the interval $J=\{x,\ldots,y\}$ because
 $\left\lceil\frac{x-1}{y-x}\right\rceil x\leq\left\lceil\frac{c-1}{\frac{c-2}{2}}\right\rceil c=\left\lceil 2+\frac{2}{c-2}\right\rceil c\leq 4c\leq 2b+c+1\leq n.$
-By Lemma 3.7 with $q=2$, $G$ has no connected partition of type $\lambda$. ∎
-Our last case to prove is Part 3, where $2c\leq b\leq\frac{c^{2}}{2}$. We first check several cases by computer.
-Proposition 3.13.
-If $2\leq c\leq 40$ and $2c\leq b\leq\frac{c^{2}}{2}$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
+By Lemma 3.9 with $q=2$, $G$ is missing a connected partition of type $\lambda$. ∎
+Our last case to prove is Part (3) of Theorem 3.2, the case where $2c\leq b\leq\frac{c^{2}}{2}$.
+Our strategy is as follows.
+Given integers $b$, $c$, and $n$ with $c\geq 2$, $2c\leq b\leq\frac{c^{2}}{2}$, and $n\geq 2b+c+1$, we wish to find a positive integer $q$ such that the integers
+(3.3) $x=\left\lceil\frac{b+1}{q}\right\rceil\text{ and }y=\left\lfloor\frac{b+c}{q}\right\rfloor$
+satisfy that $x\geq c+1\geq c_{1}+1$, $x<y$, and that there exists a partition $\lambda$ of $n$ with all parts in the interval $J=\{x,\ldots,y\}$.
+Then $G$ will have no connected partition of type $\lambda$ by Lemma 3.9.
+By Lemma 3.10, a sufficient condition to find such a $\lambda\vdash n$ is if
+(3.4) $\left\lceil\frac{x-1}{y-x}\right\rceil x\leq 2b+c+1.$
+Example 3.13.
+Let $b=75$ and $c=14$, so that $n\geq 2b+c+1=165$.
+  • $\circ$
+    For $q\geq 6$, we have $x\leq\left\lceil\frac{76}{6}\right\rceil=13\leq c$, so this will not work.
+  • $\circ$
+    For $q=5$, we have $x=\left\lceil\frac{76}{5}\right\rceil=16$ and $y=\left\lfloor\frac{89}{5}\right\rfloor=17$. We have $x\geq c+1$, but
+    $\left\lceil\frac{x-1}{y-x}\right\rceil x=240>165,$
+    so this will not work. We cannot write $n=239$ as a sum of $16$’s and $17$’s.
+  • $\circ$
+    For $q=4$, we have $x=\left\lceil\frac{76}{4}\right\rceil=19$ and $y=\left\lfloor\frac{89}{4}\right\rfloor=22$. We have $x\geq c+1$ and
+    $\left\lceil\frac{x-1}{y-x}\right\rceil x=114\leq 165,$
+    so this will work.
+Lemma 3.10 tells us that there exists a partition $\lambda$ of $n$ with all parts in the interval $J=\{19,20,21,22\}$ and Lemma 3.9 tells us that $G$ will be missing a connected partition of type $\lambda$.
+The next example shows that there may be no single value of $q$ for which the inequality (3.4) holds.
+However, for every value of $n\geq 2b+c+1$, there is some $q$ such that there exists $\lambda\vdash n$ with all parts in the interval $J=\{x,\ldots,y\}$, where $x$ and $y$ are as in (3.3).
+Example 3.14.
+Let $b=24$ and $c=7$, so that $n\geq 2b+c+1=56$.
+  • $\circ$
+    For $q\geq 4$, we have $x\leq 7\leq c$, so this will not work.
+  • $\circ$
+    For $q=3$, we have $x=9$, $y=10$, and $\left\lceil\frac{x-1}{y-x}\right\rceil x=72>56$, so this will not work.
+  • $\circ$
+    For $q=2$, we have $x=13$, $y=15$, and $\left\lceil\frac{x-1}{y-x}\right\rceil x=78>56$, so this will not work.
+  • $\circ$
+    For $q=1$, we have $x=25$, $y=31$, and $\left\lceil\frac{x-1}{y-x}\right\rceil x=100>56$, so this will not work.
+We find that no single value of $q$ satisfies (3.4).
+For $n\geq 72$, we can take $q=3$, but this will not work for every $n\geq 56$.
+Nevertheless, for each $56\leq n\leq 71$, there is some $q$ such that there exists $\lambda\vdash n$ with all parts in the interval $J=\{x,\ldots,y\}$, where $x$ and $y$ are as in (3.3).
+$q=3$ $:\{\underbrace{54,55,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}56},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}57},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}58},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}59},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}60}}_{6J}\}\cup\{\underbrace{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}63},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}64},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}65},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}66},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}67},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}68},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}69},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}70}}_{7J}\}\cup\{\underbrace{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}72},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}73},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}74},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\ldots}}_{8J}\}\cup\cdots$
+$q=2$ $:\{\underbrace{52,53,54,55,56,57,58,59,60}_{4J}\}\cup\{\underbrace{65,66,67,68,69,70,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}71},72,73,74,75}_{5J}\}\cup\cdots$
+$q=1$ $:\{\underbrace{50,51,52,53,54,55,56,57,58,59,60,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}61},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}62}}_{2J}\}\cup\{\underbrace{75,76,77,\ldots}_{3J}\}\cup\cdots$
+We will soon show by an analytic argument that if $c\geq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$, then we can find a positive integer $q$ such that the integers $x$ and $y$ as in (3.3) satisfy $x\geq c+1$, $x<y$, and the inequality (3.4).
+Our estimates will not be valid for $2\leq c\leq 499$, so we check these cases by computer.
+For $41\leq c\leq 499$, we can always find a single value of $q$ for which (3.4) holds, as in Example 3.13.
+When $2\leq c\leq 40$, we will sometimes need different values of $q$ depending on $n$ to find a suitable partition $\lambda$ of $n$, as in Example 3.14.
+Sage code used to check the following Lemma will be provided in the Appendix.
+Lemma 3.15.
+Fix positive integers $b$, $c$, and $n$ with $2\leq c\leq 499$, $2c\leq b\leq\frac{c^{2}}{2}$, and $n\geq 2b+c+1$.
+There exists a positive integer $q$ such that the integers $x$ and $y$ as in (3.3) satisfy that $x\geq c+1$, $x<y$, and there exists a partition $\lambda$ of $n$ with all parts in the interval $J=\{x,\ldots,y\}$.
 Proof.
-By Proposition 3.10, we may assume that $2b+c+1\leq n\leq\left\lceil\frac{b}{c-1}\right\rceil(b+1)$, so there are finitely many triples $(b,c,n)$ to consider.
-For each one, we checked by computer that there exists a positive integer $q$ and a partition $\lambda\vdash n$ with all parts in the corresponding interval $J=\{x,\ldots,y\}$, where $x\geq c+1\geq c_{1}+1$, so the result follows from Lemma 3.7.
+First suppose that $2\leq c\leq 40$.
+By Lemma 3.10, we can take $q=1$ if $n\geq\left\lceil\frac{b}{c-1}\right\rceil(b+1)$, so there are finitely many triples $(b,c,n)$ to consider.
+For each one, we checked by computer, using the Sage code in Method 4.1, that there exists such a positive integer $q$.
+Now suppose that $41\leq c\leq 499$.
+By Lemma 3.10, there exists such a partition $\lambda$ if the inequality (3.4) holds.
+There are finitely many pairs $(b,c)$ to consider.
+For each one, we checked by computer, using the Sage code in Method 4.2, that there exists such a positive integer $q$.
 ∎
-Lemma 3.14.
-Suppose that we can find a positive integer $q$ such that the integers $x$ and $y$ as in (3.1) satisfy $x\geq c+1$, $x<y$, and
+We now come to the most technical part of the proof, where given any $c\geq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$, we will find an integer $q$ such that the integers $x$ and $y$ as in (3.3) satisfy $x\geq c+1$, $x<y$, and the inequality
 (3.5) $\left\lceil\frac{x-1}{y-x}\right\rceil x\leq 2b+c+1.$
-Then $G$ is missing a connected partition of some type $\lambda\vdash n$.
-Proof.
-This follows directly from Lemma 3.7 and Lemma 3.8. ∎
-Proposition 3.15.
-If $41\leq c\leq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
-Proof.
-There are finitely many pairs $(b,c)$ to consider. For each one, we checked by computer that there exists a positive integer $q$ such that $x\geq c+1$, $x<y$, and (3.3) holds. ∎
-We now come to the most technical part of the proof, where given any $c\geq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$, we will find a suitable integer $q$. We first make a remark about this inequality.
+We first make a remark about this inequality.
 Remark 3.16.
@@
 To minimize this, we want $q$ to be large, and to have $x\geq c+1$, we could take $q\approx\frac{b}{c}$, which would give us $\left\lceil\frac{x-1}{y-x}\right\rceil x\approx b$, so this is promising.
-However, the floors and ceilings mean that the denominator $(y-x)$ may be closer to $(\frac{c}{q}-2)$, and this $(-2)$ term can really hurt us if $q$ is close to $\frac{c}{2}$.
+However, the floors and ceilings mean that the denominator (yx)(y-x) may be closer to $(\frac{c}{q}-2)$, and this (2)(-2) term can really hurt us if $q$ is close to $\frac{c}{2}$.
 Therefore, when $b$ is close to $\frac{c^{2}}{2}$, we will need to make a clever choice of $q$.
 Lemma 3.17.
-Fix positive integers $b$ and $c$ with $c\geq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$.
-There exists a positive integer $q$ such that, letting $x$ and $y$ be as in (3.1), we have $x\geq c+1$, $x<y$, and the inequality (3.3) holds.
+Fix positive integers $b$ and $c$ with $c\geq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$. There exists a positive integer $q$ such that, letting
+(3.6) $x=\left\lceil\frac{b+1}{q}\right\rceil\text{ and }y=\left\lfloor\frac{b+c}{q}\right\rfloor,$
+we have $x\geq c+1$, $x<y$, and the inequality
+(3.7) $\left\lceil\frac{x-1}{y-x}\right\rceil x\leq 2b+c+1.$
 Proof.
@@
 Because $b\geq\frac{c^{2}}{20}\geq 25c$, we have $q\geq 25$ and therefore
-$q=\frac{25}{26}(q+1)\geq\frac{25}{26}\frac{b}{c}.$
-Because $b\leq\frac{c^{2}}{4.2}$, we have $q\leq\frac{b}{c}\leq\frac{c}{4.2}$. Now since $y-x$ is an integer and $y-x\geq\frac{c}{q}-2\geq 2.2$, we must have $y-x\geq 3$ and
+$q\geq\frac{25}{26}(q+1)\geq\frac{25}{26}\frac{b}{c}.$
+Because $b\leq\frac{c^{2}}{4.2}$, we have $q\leq\frac{b}{c}\leq\frac{c}{4.2}$. Now since yxy-x is an integer and $y-x\geq\frac{c}{q}-2\geq 2.2$, we must have $y-x\geq 3$ and
 $y-x\geq\frac{3}{5}(y-x+2)\geq\frac{3}{5}\frac{c}{q}.$
@@
 $c\geq\sqrt{3.4}\sqrt{b}\geq\frac{\sqrt{3.4}}{0.46}q\geq 4.008q,$
-so because $y-x$ is an integer and $y-x\geq\frac{c}{q}-2>2.008$, we must have $y-x\geq 3$ and
+so because yxy-x is an integer and $y-x\geq\frac{c}{q}-2>2.008$, we must have $y-x\geq 3$ and
 $y-x\geq\frac{3}{5}(y-x+2)\geq\frac{3}{5}\frac{c}{q}.$
@@
 Let us first consider $q=q_{0}$.
-Use integer division to write $b+1=c_{0}q_{0}+r_{0}$ for some integers $c_{0}$ and $10\leq r_{0}\leq q_{0}-$ 1.
+Use integer division to write $b+1=c_{0}q_{0}+r_{0}$ for some integers $c_{0}$ and $0\leq r_{0}\leq q_{0}-1$.
 If
@@
 Now if $r_{0}<0.38q_{0}$, we will consider $q=q_{0}-1$. We have $\sqrt{3.5}q_{0}\leq\sqrt{b}\leq\sqrt{3.5}(q_{0}+1)$, so
-$3.53.5q_{0}^{2}\leq b\leq 3.5q_{0}^{2}+7q_{0}+$ 3.5 .
+$3.5q_{0}^{2}\leq b\leq 3.5q_{0}^{2}+7q_{0}+3.5.$
 Now the integer $c_{0}$ satisfies $b+1-q_{0}\leq c_{0}q_{0}\leq b+1$, so using that $q_{0}\geq 140$, we have
@@
 $b+c=b+1+c-1=(c_{0}+3)(q_{0}-1)+c+r-1>(c_{0}+6)(q_{0}-1),$
-so at least three multiples of $(q_{0}-1)$, namely $(c_{0}+4)(q_{0}-1)$, $(c_{0}+5)(q_{0}-1)$, and $(c_{0}+6)(q_{0}-1)$, lie strictly between $(b+1)$ and $(b+c)$.
+so at least three multiples of $(q_{0}-1)$, namely $(c_{0}+4)(q_{0}-1)$, $(c_{0}+5)(q_{0}-1)$, and $(c_{0}+6)(q_{0}-1)$, lie strictly between $b+1$ and $b+c$.
 This means that $x\leq c_{0}+4$ and $y\geq c_{0}+6$, so $y-x\geq 2$.
@@
 This completes the proof. ∎
-Now we can prove the remaining Part 3 of Theorem 3.1.
-Proposition 3.18.
-If $c\geq 2$ and $2c\leq b\leq\frac{c^{2}}{2}$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
-Proof.
-This follows directly from Proposition 3.13, Proposition 3.15, Lemma 3.14, and Lemma 3.17. ∎
+Now we can prove the remaining Part (3) of Theorem 3.2, the case where $2c\leq b\leq\frac{c^{2}}{2}$.
+Proof of Part (3) of Theorem 3.2. .
+If $2\leq c\leq 499$, then Lemma 3.15 tells us that there is an integer $q$ and a partition $\lambda$ of $n$ with all parts in the corresponding interval $J=\{x,\ldots,y\}$, where $x$ and $y$ are as in (3.3).
+If $c\geq 500$, then Lemma 3.17 and Lemma 3.10 again give us these $q$ and $\lambda\vdash n$.
+Either way, now Lemma 3.9 tells us that $G$ is missing a connected partition of this type $\lambda\vdash n$.
+∎
 We also prove that spiders with four legs are not $e$-positive.
 Proof of Theorem 1.6. .
-Because $S$ has four legs, we have $c\geq c_{1}+1\geq 2$. If $b\leq\frac{c^{2}}{2}$, the result follows from Parts 1, 2, and 3 of Theorem 3.1. If $b\geq\frac{c^{2}}{2}$, then
+Let $S$ be a spider with four legs, so that $c\geq c_{1}+1\geq 2$. If $b\leq\frac{c^{2}}{2}$, then the result follows from Parts (1), (2), and (3) of Theorem 3.2. If $b\geq\frac{c^{2}}{2}$, then
 $n\geq 2b+c+1\geq c^{2}+c+1,$
-and Zheng showed [spiderepos, Corollary 4.6] that such a spider $S$ is not $e$-positive. ∎
-In order to prove Conjecture 1.1, it now remains to consider trees with a vertex of degree $4$ that is adjacent to a leaf, such as the spider $S(6,4,1,1)$.
+and Zheng showed [19, Corollary 4.6] that such a spider $S$ is not $e$-positive. ∎
+In order to prove Conjecture 1.2, it now remains to consider trees with a vertex of degree $4$ that is adjacent to a leaf, such as the spider $S(6,4,1,1)$ from Figure 1.
 It turns out that $S(6,4,1,1)$ has a connected partition of every type $\lambda\vdash 13$, which shows that Wolfgang’s necessary condition will not be enough.
 We conclude by showing that there are infinitely many such examples.
-Figure 3. The spider $S(6,4,1,1)$ has a connected partition of every type $\lambda\vdash 13$
-$S(6,4,1,1)$
 Proposition 3.18.
-Let $m\geq 1$ be an integer. The spider $S(6m,6m-2,1,1)$ has a connected partition of type $\lambda$ for every $\lambda\vdash 12m+1$.
+Let $m\geq 1$ be an integer. The spider S(6m,6m2,1,1)S(6m,6m-2,1,1) has a connected partition of type $\lambda$ for every $\lambda\vdash 12m+1$.
+Note that by Theorem 1.6, the spider S(6m,6m2,1,1)S(6m,6m-2,1,1) is not $e$-positive.
 Proof.
-We claim that if $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$ has a rearrangement $\alpha$ with $6m-1,6m\notin\text{set}(\alpha)$, then $S(6m,6m-2,1,1)$ has a connected partition of type $\lambda$.
-Label the vertices on the leg of length $26m-$ 2 as $u_{1},\ldots,u_{6m-2}$, moving in from the leaf, and label the vertices on the leg of length $6m$ as $v_{1},\ldots,v_{6m}$ similarly.
+We claim that if $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$ has a rearrangement $\alpha$ with $6m-1,6m\notin\text{set}(\alpha)$, then S(6m,6m2,1,1)S(6m,6m-2,1,1) has a connected partition of type $\lambda$.
+Label the vertices on the leg of length 6m26m-2 as $u_{1},\ldots,u_{6m-2}$, moving in from the leaf, and label the vertices on the leg of length $6m$ as $v_{1},\ldots,v_{6m}$ similarly.
 Suppose that $\alpha$ is a rearrangement of $\lambda$ that satisfies
-$\alpha_{1}+\cdots+\alpha_{i}\leq 6m-2\text{ and }\alpha_{1}+\cdots+\alpha_{i+1}\geq 6m+1.$
-Then we can partition the vertices $\{u_{1},u_{2},\ldots,u_{\alpha_{1}+\cdots+\alpha_{i}}\}$ into paths of sizes $\alpha_{1},\ldots,\alpha_{i}$, we can partition the vertices $\{v_{1},v_{2},\ldots,v_{\alpha_{i+2}+\cdots+\alpha_{\ell}}\}$ into paths of sizes $\alpha_{i+2},\ldots,\alpha_{\ell}$, and the remaining vertices form a connected set of size $\alpha_{i+1}$.
+[…]
+Then we can partition the vertices $\{u_{1},u_{2},\ldots,u_{\alpha_{1}+\cdots+\alpha_{i}}\}$ into paths of orders $\alpha_{1},\ldots,\alpha_{i}$, we can partition the vertices $\{v_{1},v_{2},\ldots,v_{\alpha_{i+2}+\cdots+\alpha_{\ell}}\}$ into paths of orders $\alpha_{i+2},\ldots,\alpha_{\ell}$, and the remaining vertices form a connected set of order $\alpha_{i+1}$.
 This establishes our claim.
-Note that if $\lambda$ has at least two $1$’s or if $\lambda=(2,2,\ldots,2,1)$, then $S(6m,6m-2,1,1)$ has a connected partition of type $\lambda$, so we may assume that $\lambda$ has at most one $1$ and has a part $x\geq 3$.
+Note that if $\lambda$ has at least two $1$’s or if $\lambda=(2,2,\ldots,2,1)$, then S(6m,6m2,1,1)S(6m,6m-2,1,1) has a connected partition of type $\lambda$, so we may assume that $\lambda$ has at most one $1$ and has a part $x\geq 3$.
 By our claim, it suffices to find a rearrangement $\alpha$ of $\lambda$ with $6m-1,6m\notin\text{set}(\alpha)$.
 Fix a rearrangement $\beta$ of $\lambda$.
-We may assume that $16m-$ 1 or $6m\in\text{set}(\beta)$, otherwise we could take $\alpha=\beta$.
-Letting $\beta^{*}=(\beta_{\ell},\ldots,\beta_{1})$ denote the reverse composition, we may assume that $16m-$ 1 or $6m\in\text{set}(\beta^{*})$, otherwise we could take $\alpha=\beta^{*}$.
+We may assume that 6m16m-1 or $6m\in\text{set}(\beta)$, otherwise we could take $\alpha=\beta$.
+Letting $\beta^{*}=(\beta_{\ell},\ldots,\beta_{1})$ denote the reverse composition, we may assume that 6m16m-1 or $6m\in\text{set}(\beta^{*})$, otherwise we could take $\alpha=\beta^{*}$.
 Now there are only a few cases to consider.
@@
 has $\text{set}(\alpha)=\{\ldots,6m-x-1,6m-x+1,6m+1,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$.
-Otherwise, the parts $(\beta_{1},\ldots,\beta_{i})$ consist of a single $1$ and $(3m-1)$ $2$’s, some $\beta_{k}=x\geq 3$ for $i+2\leq k\leq\ell$, and the composition
+Otherwise, the parts $(\beta_{1},\ldots,\beta_{i})$ consist of a single $1$ and (3m1)(3m-1) 22’s, some $\beta_{k}=x\geq 3$ for $i+2\leq k\leq\ell$, and the composition
 $\alpha=(2,2,\ldots,2,2,x,1,2,\beta_{i+2},\ldots,\beta_{k-1},\beta_{k+1},\ldots,\beta_{\ell})$
@@
 Now we may assume that every $1\leq k\leq i$ has $\beta_{k}=2$ or $3$, and by reversing, that every $i+2\leq k\leq\ell$ also has $\beta_{k}=2$ or $3$.
-Because $16m-$ 1 is not divisible by either $2$ or $3$, there must be some $1\leq k\leq i$ with $\beta_{k}=3$ and some $i+2\leq k^{\prime}\leq\ell$ with $\beta_{k^{\prime}}=2$.
+Because 6m16m-1 is not divisible by either $2$ or $3$, there must be some $1\leq k\leq i$ with $\beta_{k}=3$ and some $i+2\leq k^{\prime}\leq\ell$ with $\beta_{k^{\prime}}=2$.
 Now the composition
@@
 has $\text{set}(\alpha)=\{\ldots,6m-4,6m-2,6m+1,6m+4,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$. This completes the proof. ∎
+4. Appendix
+Method 4.1.
+This is used to check that for integers $b$, $c$, and $n$ with $2\leq c\leq 40$, $2c\leq b\leq\frac{c^{2}}{2}$, and $2b+c+1\leq n\leq\left\lceil\frac{b}{c-1}\right\rceil(b+1)$, we can find a positive integer $q$ such that the integers $x$ and $y$ as in (3.3) satisfy $x\geq c+1$, $x<y$, and $tx\leq n\leq ty$ for some $t$; and therefore that there exists a partition $\lambda$ of $n$ with all parts in the interval $J=\{x,\ldots,y\}$.
+1def checkc40():
+2 for c in range(2,41):
+3 for b in range(2*c,floor(c^2/2)+1):
+4 for n in range(2*b+c+1,ceil(b/(c-1))*(b+1)+1):
+5 print(’b:’,b,’c:’,c,’n:’,n)
+6 foundq=false
+7 for q in range(1,ceil((b+1)/c)):
+8 x=ceil((b+1)/q)
+9 y=floor((b+c)/q)
+10 t=floor(n/x)
+11 if x>=c+1 and x<y and t*x<=n<=t*y:
+12 print(’q:’,q,’x:’,x,’y:’,y,’t:’,t)
+13 foundq=true
+14 if not foundq:
+15 return false
+16 return true
+Method 4.2.
+This is used to check that for integers $b$ and $c$ with $2\leq c\leq 40$ and $2c\leq b\leq\frac{c^{2}}{2}$, we can find a positive integer $q$ such that the integers $x$ and $y$ as in (3.3) satisfy $x\geq c+1$, $x<y$, and the inequality (3.4); and therefore that there exists a partition $\lambda$ of $n$ with all parts in the interval $J=\{x,\ldots,y\}$.
+1def checkc500():
+2 for c in range(41,500):
+3 for b in range(2*c,floor(c^2/2)+1):
+4 print(’b:’,b,’c:’,c)
+5 foundq=false
+6 for q in range(1,ceil((b+1)/c)):
+7 x=ceil((b+1)/q)
+8 y=floor((b+c)/q)
+9 M=ceil((x-1)/(y-x))*x
+10 if x>=c+1 and x<y and M<=2*b+c+1:
+11 print(’q:’,q,’x:’,x,’y:’,y,’M:’,M)
+12 foundq=true
+13 if not foundq:
+14 return false
+15 return true
 5. Acknowledgments
-The author would like to think José Aliste-Prieto, Richard Stanley, Stephanie van Willigenburg, and Kai Zheng for helpful discussions.
-[…]
+The author would like to thank José Aliste-Prieto, Richard Stanley, Stephanie van Willigenburg, and Kai Zheng for helpful discussions.